Use the following to answer the next three questions s lesa than the national average of 0.124. Data were collected from a random sample of 547 women and among a) He p-0.124, H:p>0.124 b) Hk p 0082; H, p 0082 0082-0.124 0.124-0082 0.124-0.082 32) The p-value for this test is 0.0014. Based on this make? a) Fail to reject H. There is not sufficient evidence to conclude that the proportion of women diagnosed with breast cancer is less than 0.124 Reject Ho and conclude that the proportion of women diagnosed with breast cancer is less than 0.124. b) to 0.124. 33) test on the b) 09463
Answer #1

30. Here claim is that p<0.124

So hypothesis is d. H0 : p 0.124 vs H_a:p<0.124

31. Test statistics is

z=\frac{\widehat{p}-p}{\sqrt{\frac{pq}{n}}}=\frac{0.082-0.124}{\sqrt{\frac{0.124(1-0.124)}{547}}}

So b is correct answer.

32. As P value is less than alpha, we reject the null hypothesis which means we have sufficient evidence to support the claim that p<0.124

So b is correct answer

33. As it is right tailed test and test statistics is z=1.61

So P value is P(z>1.61)=0.0537

So d is correct answer